can you solve this system of linear equation by graphing please and show me how you did it.
Like how you found the equations to make the graph
353 people attend a local play at a private club. Members get tickets for $2.75, while non-members have to pay $6.50. If the total gate for the play came to $1762, how many members and non-members came to the play?

Answer by richard1234(7193)

Let m be the number of members and n be the number of non-members. We know that 353 people showed up, so m+n = 353. Also, each member paid $2.75 and each non-member paid $6.50, so we have 2.75m + 6.50n = 1762. So we have the following system of linear equations:

m+n = 353
2.75m + 6.50n = 1762

Then you can graph both equations in the mn-plane. However the best way to solve it is likely to use substitution (n = 353-m, substitute).